# Non-Markovian character and irreversibility of real-time quantum many-body dynamics

### v1

@@ VI.2 Non-Markovian character of fission dynamics
In Figs. 2 and 3 we plot the sum of the absolute changes in the single-particle occupation probabilities at some fixed time intervals fm/c and the absolute differences between the initial and time dependent single-particle occupation probabilities. Similar results have been reported in Ref. Bulgac (2022), however, with slightly different initial conditions for the same nucleus U with the same NEDF SeaLL1, but for smaller values of fm/c. In that case the initial quasiparticle wave functions were obtained using the code HFBTHO Navarro Perez et al. (2017), placed on the 3D spatial lattice, and only adjusting the proton and neutron chemical potentials to fix the correct average particle numbers. In all the simulations reported here we have run the static SLDA code on the 3D spatial lattice until full self-consistency was achieved. Since the phase space is much larger on a 3D spatial lattice than that used in HFBTHO code and since the kinetic energy and anomalous densities are formally diverging in a 3D space Bulgac (1980), the self-consistent SLDA equations need to be regularized and renormalized Bulgac and Yu (2002); Jin et al. (202). One important benefit of performing this additional self-consistency evaluation of the SLDA equations on the 3D spatial lattice is a much more numerically accurate solution of the TDSLDA equations.
@@
The quantum mechanical nature of the nuclear shape proves however to be more complex than envisioned by Bertsch Bertsch (1980); Bertsch et al. (1987); Barranco et al. (1990, 1988); Bertsch and Bulgac (1997); Bulgac (2022) and these jumps appear to be highly correlated in time, which is a qualitatively new aspect of fission dynamics in particular. This aspect is particularly interesting, since as we have proven in earlier fission simulations Bulgac et al. (2016b, 2019, 2020), the descent from the top of the outer barrier to the scission configuration is a highly dissipative process, in which case one would expect that stochasticity of the dynamics might play a crucial role. In the presence of a strong dissipation, stochasticity in case of classical dynamics and in numerous phenomenological fission models is modelled with a Langevin force Vogt et al. (2009); Vogt and Randrup (2013); Verbeke et al. (2018); Albertsson et al. ; Albertsson et al. (2020); Ishizuka et al. (2017); Randrup and Möller (2011); Randrup et al. (2011); Randrup and Möller (2013); Sierk (2017); Möller et al. (2001, 2004); Litaize et al. (2012); Becker et al. (2013). This is a qualitatively new situation in non-equilibrium dynamics, so far never discussed in literature as far as we can judge, where in the presence of strong dissipation, memory effects are also very strong and the single-particle occupation probabilities dynamics show a clear non-Markovian behavior.
@@
The quantity appears to change at two different rates for times smaller than 500-600 fm/c and at a slower rather for larger times. We see three different sources for this behavior. i) With time the strength of the pairing correlations and the absolute magnitude of the pairing gaps decreases, though it does never vanish, see Fig. 4 and Bulgac et al. (2016b, 2019, 2020). ii) The fissioning nucleus and the FFs after separation still convert deformation energy into thermal energy, which leads to increasing occupation probabilities of higher energy single-particle states. iii) The neck appears to start forming at times 500-700 fm/c Bulgac et al. (2019); Bulgac (2022); Bulgac et al. (2023) and scission occurs at time 2,300-2,700 fm/c, depending on the initial conditions considered. In all TDDFT simulations one has to choose the initial nuclear shape but imposing a shape constraint and requiring that the total energy is near the outer fission barrier. The shape constraint is not inherent in the many-body Schrödinger equation and as in the initial Bohr and Wheeler (1939) paper this is merely a theoretical tool used since 1939. The nucleus during its descent from the top of the fission barrier needs to adjust to the absence of the artificial shape constraint imposed on the initial state. For times larger than 500-600 fm/c the nucleus appears to have settled to a different slower, but rather well defined rate, a bit higher for the neutron system than for the proton system. The neck, depending on its size, increasingly impedes matter, linear and angular momentum, and energy exchange between the two emerging FFs, and that is another reason why the rate of “equilibration” we see in Fig. 3 settles to a smaller value. At the same time, scission, which occurs at much later times, between 2,300-2,700 fm/c depending on the initial conditions and NEDF used, does not appear to affect this rate.
@@
VI.3 Particle projection and use of a reduced set of canonical wave functions in time-dependent simulations
We have estimated average discrepancies between the particle projected and unprojected number densities at various times
@@
In Fig. 6 we show the total energy U as a function of time, depending on various numbers of canonical wave functions used as an initial set. With the exception of the simulations with the entire set of quasi-particle wave functions and the entire set of canonical wave functions used as initial conditions, which are indistinguishable on this plot, every run performed with any number of wave functions as large as 50,000 initial canonical wave functions, failed to lead to fission, and moreover the total energy of the system is not conserved, the of the entire system basically does not change in time. and the nucleus only keeps “heating up.” For any set of initial conditions, using the canonical wave functions with largest occupation probabilities in the interval , the nucleus does not fission, but only heats up. In the lower panel we display the behavior of the quadruple moment of the entire nucleus as a function of time for the same choice of initial conditions as for the upper panel. Only for the case when the entire set of quasi-particle or canonical wave functions is used the nucleus fissions, otherwise its size remains basically unchanged as a function of time. In this case the initial condition had an additional excitation energy of about 1.17 MeV and the simulations performed with a reduced set of canonical wave function still did not fission.
@@
The vast majority of simulations performed by other authors use the TDHF+BCS or TDHF+TDBCS approximation Zhang et al. (2023); Li et al. (2023a, b); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps and Hashimoto (2019), where TD stands for time-dependent and HF for Hartree-Fock. Either the BCS or TDBCS approximations are further approximations to the TD Hartree-Fock-Bogoliubov (TDHFB) equations with cutoffs in the number of levels allowed to participate in pairing. Moreover, both BCS and TDBCS assume the spatial profiles of the and -components of the quasi-particle wave functions are identical, as in the initial BCS approximation Bardeen et al. (1957) for weak pairing correlations. Even further, the continuity equation is violated in the TDHF+TDBCS approximation Scamps et al. (2012), which is still widely used today despite this Zhang et al. (2023); Li et al. (2023a, b); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps and Hashimoto (2019) (to list a few studies).
Both BCS and TDBCS represent more limited approximations than evolving a truncated set of canonical quasi-particle wave functions, or using a spherical cutoff, as was originally done in Bulgac et al. (2016b). As described above, all trajectories performed in the canonical basis with a cutoff didn’t fission. This is likely the reason why BCS and TDBCS simulations of (induced) fission Zhang et al. (2023); Li et al. (2023a, b); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps and Hashimoto (2019) start with initial compound nuclei on the potential energy surface of the fissioning nucleus that are well below the saddle point state of the nucleus considered. Only in such situations, for configurations where the Coulomb repulsion considerably exceeds the nuclear surface energy, could such simulations produce separated fission fragments.
@@
The projected particle probability distributions , see Eq. (130), for protons and neutrons obtained are shown in Fig. 7, are slightly asymmetric with respect to the average proton and neutron numbers. In Figs. 8-10 we show the differences between the unprojected and particle-projected number densities for protons and neutrons in the case of the induced fission of U at the top of the outer barrier, at the scission configurations, and for fully separated FFs. The densities were evaluated with 5,000 canonical quasi-particle wave functions, in which case the agreement at each time during the evolution with the entire set had very small relative errors, see also Fig 5.
@@ VI.4 Irreversibility in isolated quantum systems
Another relevant aspect for the nuclear dynamics, which can be revealed with the help of canonical wave functions, is the irreversible time evolution of an isolated nuclear system, before it emits any nucleons or before it couples with the electromagnetic fields and emits photons or later on -particles after coupling with the weak interactions. An isolated excited nucleus has a vanishing von Neumann or Shannon entropy, which naively would point to the absence of any irreversible time-evolution of such a system, which clearly is incorrect. At the classical level one can evaluate the Boltzmann entropy of an excited system, which would clearly characterize the irreversible time evolution of the system. The non-equilibrium evolution of isolated quantum systems can be characterized however with the help of the entanglement entropy Calabrese and Cardy (2005, 2006); Alba and Calabrese (2017). The entanglement entropy is non-vanishing even in the ground states of interacting systems Srednicki (1993). There is however no unique definition of the entanglement entropy, as it is well known Nordheim (1928); Uehling and Uhlenbeck (1933); Srednicki (1993); Klich (2006); Boguslawski and Tecmer (2014); Klich (2006); Amico et al. (2008); Horodecki et al. (2009); Haque et al. (2009); Eisert et al. (2010); Boguslawski and Tecmer (2014); Gigena and Rossignoli (2015); Bengtsson and Życzkowski (2017); Johnson (2018); Robin et al. (2021); Hoppe et al. (2021); Johnson and Gorton (2022); Bulgac (2022, 2023a, 2023b); Tichai et al. ; Fasano et al. (2022); Bulgac et al. (2023); Robin and Savage (2023); Gu et al. (2023); Pandharipande et al. (1984); Stoitsov et al. (1993); Reinhard et al. (1999); Dobaczewski et al. (1996); Tajima (2004); Tichai et al. (2019); Fasano et al. (2022); Chen et al. (2022); Hu et al. (2022); Hagen et al. (2022); Kortelainen et al. (2022). For nuclear systems, and particularly for heavy nuclei, which have an enormous number of degrees of freedom only the orbital entropy can be evaluated in the near future
@@
Here we will illustrate the irreversible fission dynamics in the largest simulation we have performed on a spatial lattice , which required 4,609 nodes with 27,654 GPUs on the Summit supercomputer for about 15 wall-hours for a single fission trajectory, in one of the largest (if not the largest) direct numerical simulation ever reported, see Fig. 11. Similar results have been reported in Ref. Bulgac (2022, 2023a, 2023b); Bulgac et al. (2023) for both fission and for a U+U collision at 1,500 MeV in center-of-mass frame, however for much smaller simulation boxes. The orbital entanglement entropy is very large initially, as expected in a system with very strong pairing correlations. As the nucleus evolves towards scission it heats up, achieving temperatures above 1 MeV Bulgac et al. (2016b, 2019, 2020) and the pairing correlations weaken, but they do not disappear, see Refs. Bulgac et al. (2016b, 2019, 2020) and Fig. 4. As Magierski et al. (2022) have recently shown, even in the high energy collisions of Zr+Zr, due to the excitation of the Higgs pairing mode Barankov and Levitov (2006); Bulgac and Yoon (2009), pairing correlations, which are absent in the initial nuclei, acquire a large amplitude at very large excitation energies in the proximity of the Coulomb barrier of the colliding nuclei of the compound nucleus formed in this collision. While approaching the scission configuration, the matter exchange between the two halves of the fissioning nucleus slows down and completely stops after scission, which in the Fig. 11 is around 1,000 fm/c. After the two FFs separate, they are highly excited, and in particular the light FF is also very highly deformed, and both fragment relax, and the entanglement entropy increases with time, as the single-particle levels are repopulated, reaching values almost equal to the initial entanglement entropy of the cold strongly correlated nucleus near the top of the outer barrier. The temperature of the final FFs is significantly larger, the remaining pairing correlations are weaker, and while one process favors more occupation particle redistribution, the other works in the opposite direction.
@@
In the large simulation box we studied the fission of U we thus have clearly identified two regimes. In this very large simulation box the FFs still evolve in time when they reach the box walls and they clearly did not reach thermalization. One might be tempted to isolate each FF separately and follow its evolution in its own center-of-mass. This can appear physically motivated, as at sufficiently large spatial separations one expects that the FFs can hardly influence each other any longer, apart from the relatively weak Coulomb field. It is however not clear how one can formally proceed, since apart from the quasiparticle wave functions, which can be very well localized inside a specific FF, the time evolution of a specific FF is also controlled by the component of the quasiparticle wave functions, which are mostly fully delocalized, as we discussed in Section III, and through these components the two FFs can “communicate” with each other. This formal aspect of the TDDFT has not been developed yet.

### v2

@@ VI.2 Non-Markovian character of fission dynamics
In Figs. 2 and 3 we plot the sum of the absolute changes in the single-particle occupation probabilities at some fixed time intervals fm/c and the absolute differences between the initial and time dependent single-particle occupation probabilities. Similar results have been reported in Ref. Bulgac (2022), however, with slightly different initial conditions for the same nucleus 236U with the same NEDF SeaLL1, but for smaller values of fm/c. In that case the initial quasiparticle wave functions were obtained using the code HFBTHO Navarro Perez et al. (2017), placed on the 3D spatial lattice, and only adjusting the proton and neutron chemical potentials to fix the correct average particle numbers. In all the simulations reported here we have run the static SLDA code on the 3D spatial lattice until full self-consistency was achieved. Since the phase space is much larger on a 3D spatial lattice than that used in HFBTHO code and since the kinetic energy and anomalous densities are formally diverging in a 3D space Bulgac (1980), the self-consistent SLDA equations need to be regularized and renormalized Bulgac and Yu (2002); Jin et al. (202). One important benefit of performing this additional self-consistency evaluation of the SLDA equations on the 3D spatial lattice is a much more numerically accurate solution of the TDSLDA equations.
@@
The quantum mechanical nature of the nuclear shape proves however to be more complex than envisioned by Bertsch Bertsch (1980); Bertsch et al. (1987); Barranco et al. (1990, 1988); Bertsch and Bulgac (1997); Bulgac (2022) and these jumps appear to be highly correlated in time, which is a qualitatively new aspect of fission dynamics in particular. This aspect is particularly interesting, since as we have proven in earlier fission simulations Bulgac et al. (2016b, 2019, 2020), the descent from the top of the outer barrier to the scission configuration is a highly dissipative process, in which case one would expect that stochasticity of the dynamics might play a crucial role. In the presence of a strong dissipation, stochasticity in case of classical dynamics and in numerous phenomenological fission models is modelled with a Langevin force Vogt et al. (2009); Vogt and Randrup (2013); Verbeke et al. (2018); Albertsson et al. (2020); Ishizuka et al. (2017); Randrup and Möller (2011); Randrup et al. (2011); Randrup and Möller (2013); Sierk (2017); Möller et al. (2001, 2004); Litaize et al. (2012); Becker et al. (2013). This is a qualitatively new situation in non-equilibrium dynamics, so far never discussed in literature as far as we can judge, where in the presence of strong dissipation, memory effects are also very strong and the single-particle occupation probabilities dynamics show a clear non-Markovian behavior.
@@
The quantity appears to change at two different rates for times smaller than 500-600 fm/c and at a slower rather for larger times. We see three different sources for this behavior. i) With time the strength of the pairing correlations and the absolute magnitude of the pairing gaps decreases, though it does never vanish, see Fig. 4 and Bulgac et al. (2016b, 2019, 2020). ii) The fissioning nucleus and the FFs after separation still convert deformation energy into thermal energy, which leads to increasing occupation probabilities of higher energy single-particle states. iii) The neck appears to start forming at times 500-700 fm/c Bulgac et al. (2019); Bulgac (2022); Bulgac et al. (2023) and scission occurs at time 2 300-2 700 fm/c, depending on the initial conditions considered. In all TDDFT simulations one has to choose the initial nuclear shape but imposing a shape constraint and requiring that the total energy is near the outer fission barrier. The shape constraint is not inherent in the many-body Schrödinger equation and as in the initial Bohr and Wheeler (1939) paper this is merely a theoretical tool used since 1939. The nucleus during its descent from the top of the fission barrier needs to adjust to the absence of the artificial shape constraint imposed on the initial state. For times larger than 500-600 fm/c the nucleus appears to have settled to a different slower, but rather well defined rate, a bit higher for the neutron system than for the proton system. The neck, depending on its size, increasingly impedes matter, linear and angular momentum, and energy exchange between the two emerging FFs, and that is another reason why the rate of “equilibration” we see in Fig. 3 settles to a smaller value. At the same time, scission, which occurs at much later times, between 2 300-2 700 fm/c depending on the initial conditions and NEDF used, does not appear to affect this rate.
@@
VI.3 Particle-number projection and use of a reduced set of canonical wave functions in time-dependent simulations
We have estimated average discrepancies between the particle-number projected and unprojected number densities obtained within TDSLDA at various times
@@
In Fig. 6 we show the total energy 236U as a function of time, depending on various numbers of canonical wave functions used as an initial set. With the exception of the simulations with the entire set of quasi-particle wave functions and the entire set of canonical wave functions used as initial conditions, which are indistinguishable on this plot, every run performed with any number of wave functions as large as 50 000 initial canonical wave functions, failed to lead to fission, and moreover the total energy of the system is not conserved, the of the entire system basically does not change in time, and the nucleus only keeps “heating up.” For any set of initial conditions, using the canonical wave functions with largest occupation probabilities in the interval , the nucleus does not fission, but only heats up. In the lower panel we display the behavior of the quadruple moment of the entire nucleus as a function of time for the same choice of initial conditions as for the upper panel. Only for the case when the entire set of quasi-particle or canonical wave functions is used the nucleus fissions, otherwise its size remains basically unchanged as a function of time. In this case the initial condition had an additional excitation energy of about 1.17 MeV and the simulations performed with a reduced set of canonical wave function still did not fission.
@@
The vast majority of simulations performed by other authors use the TDHF+BCS or TDHF+TDBCS approximation Zhang et al. (2024); Li et al. (2024, 2023); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps et al. (2012), where TD stands for time-dependent and HF for Hartree-Fock. Either the BCS or TDBCS approximations are further approximations to the TD Hartree-Fock-Bogoliubov (TDHFB) equations with cutoffs in the number of levels allowed to participate in pairing. Moreover, both BCS and TDBCS assume the spatial profiles of the and -components of the quasi-particle wave functions are identical, as in the initial BCS approximation Bardeen et al. (1957) for weak pairing correlations. Even further, the continuity equation is violated in the TDHF+TDBCS approximation Scamps et al. (2012), which is still widely used today despite this Zhang et al. (2024); Li et al. (2024, 2023); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps et al. (2012) (to list a few studies). In Ref. Scamps and Hashimoto (2019) a TDHFB implementation in a reduced basis set was used, which according to our present analysis might be problematic.
Both BCS and TDBCS represent more limited approximations than evolving a truncated set of canonical quasi-particle wave functions, or using a spherical cutoff, as was originally done in Bulgac et al. (2016b). As described above, all trajectories performed in the canonical basis with a cutoff didn’t fission. This is likely the reason why BCS and TDBCS simulations of (induced) fission Zhang et al. (2024); Li et al. (2024, 2023); Ren et al. (2022a, b); Bender and et al (2020); Scamps et al. (2015); Scamps and Simenel (2018); Scamps et al. (2012) start with initial compound nuclei on the potential energy surface of the fissioning nucleus that are well below the saddle point state of the nucleus considered. Only in such situations, for configurations where the Coulomb repulsion considerably exceeds the nuclear surface energy, could such simulations produce separated fission fragments.
@@
The projected particle number probability distributions , see Eq. (79), for protons and neutrons obtained are shown in Fig. 7, are slightly asymmetric with respect to the average proton and neutron numbers. In Figs. 8-10 we show the differences between the unprojected and particle-number-projected number densities for protons and neutrons in the case of the induced fission of 236U at the top of the outer barrier, at the scission configurations, and for fully separated FFs. From the time-dependent full set of QPWFs we evaluated particle-number projected densities with only 5 000 canonical quasi-particle wave functions. The agreement between the particle-number projected and unprojected densities at each selected time during the evolution had very small relative errors, see also Fig. 5.
@@ VI.4 Irreversibility in isolated quantum systems
Another relevant aspect for the nuclear dynamics, which can be revealed with the help of canonical wave functions, is the irreversible time evolution of an isolated nuclear system, before it emits any nucleons or before it couples with the electromagnetic fields and emits photons or later on -particles after coupling with the weak interactions. An isolated excited nucleus has a vanishing von Neumann or Shannon entropy, which naively would point to the absence of any irreversible time-evolution of such a system, which clearly is incorrect. At the classical level one can evaluate the Boltzmann entropy of an excited system, which would clearly characterize the irreversible time evolution of the system. The non-equilibrium evolution of isolated quantum systems can be characterized however with the help of the entanglement entropy Calabrese and Cardy (2005, 2006); Alba and Calabrese (2017). The entanglement entropy is non-vanishing even in the ground states of interacting systems Srednicki (1993). There is however no unique definition of the entanglement entropy, as it is well known Nordheim (1928); Uehling and Uhlenbeck (1933); Srednicki (1993); Klich (2006); Boguslawski and Tecmer (2014); Klich (2006); Amico et al. (2008); Horodecki et al. (2009); Haque et al. (2009); Eisert et al. (2010); Boguslawski and Tecmer (2014); Gigena and Rossignoli (2015); Bengtsson and Życzkowski (2017); Johnson (2018); Robin et al. (2021); Hoppe et al. (2021); Johnson and Gorton (2023); Bulgac (2022, 2023a, 2023b); Tichai et al. (2023); Fasano et al. (2022); Bulgac et al. (2023); Robin and Savage (2023); Gu et al. (2023); Pandharipande et al. (1984); Stoitsov et al. (1993); Reinhard et al. (1999); Dobaczewski et al. (1996); Tajima (2004); Tichai et al. (2019); Fasano et al. (2022); Chen et al. (2022); Hu et al. (2022); Hagen et al. (2022); Kortelainen et al. (2022). For nuclear systems, and particularly for heavy nuclei, which have an enormous number of degrees of freedom only the orbital entropy can be evaluated in the near future
@@
Here we will illustrate the irreversible fission dynamics in the largest simulation we have performed on a spatial lattice , which required 4 609 nodes with 27 654 GPUs on the Summit supercomputer for about 15 wall-hours for a single fission trajectory, in one of the largest (if not the largest) direct numerical simulation ever reported, see Fig. 11. Similar results have been reported in Ref. Bulgac (2022, 2023a, 2023b); Bulgac et al. (2023) for both fission and for a 238U+238U collision at 1 500 MeV in center-of-mass frame, however for much smaller simulation boxes. The orbital entanglement entropy is very large initially, as expected in a system with very strong pairing correlations. As the nucleus evolves towards scission it heats up, achieving temperatures above 1 MeV Bulgac et al. (2016b, 2019, 2020) and the pairing correlations weaken, but they do not disappear, see Refs. Bulgac et al. (2016b, 2019, 2020) and Fig. 4. As Magierski et al. (2022) have recently shown, even in the high energy collisions of 90Zr+90Zr, due to the excitation of the Higgs pairing mode Barankov and Levitov (2006); Bulgac and Yoon (2009), pairing correlations, which are absent in the initial nuclei, acquire a large amplitude at very large excitation energies in the proximity of the Coulomb barrier of the colliding nuclei of the compound nucleus formed in this collision. While approaching the scission configuration, the matter exchange between the two halves of the fissioning nucleus slows down and completely stops after scission, which in the Fig. 11 is around 1 000 fm/c. After the two FFs separate, they are highly excited, and in particular the light FF is also very highly deformed, and both fragment relax, and the entanglement entropy increases with time, as the single-particle levels are repopulated, reaching values almost equal to the initial entanglement entropy of the cold strongly correlated nucleus near the top of the outer barrier. The temperature of the final FFs is significantly larger, the remaining pairing correlations are weaker, and while one process favors more occupation particle redistribution, the other works in the opposite direction.
@@
In the large simulation box we studied the fission of 236U we thus have clearly identified two regimes. In this very large simulation box the FFs still evolve in time when they reach the box walls and they clearly did not reach thermalization. One might be tempted to isolate each FF separately and follow its evolution in its own center-of-mass. This can appear physically motivated, as at sufficiently large spatial separations one expects that the FFs can hardly influence each other any longer, apart from the relatively weak Coulomb field. It is however not clear how one can formally proceed, since apart from the quasiparticle wave functions, which can be very well localized inside a specific FF, the time evolution of a specific FF is also controlled by the component of the quasiparticle wave functions, which are mostly fully delocalized, as we discussed in Section III, and through these components the two FFs can “communicate” with each other. This formal aspect of the TDDFT has not been developed yet.